HUBBLE SPACE TELESCOPE DAILY REPORT #917 PERIOD COVERED: 0800 EST Daily Status Report as of 179/1200Z. 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed FOS/BL 3483 (Lyman Alpha Emission in Galaxies: the Case of Galaxies Causing the Z = 0) The Faint Object Spectrograph was used to make Lyman Alpha emission line observations of the galaxy PKS-1127-1 which is near the line of sight of a known quasi-stellar object. This galaxy is expected to produce Magnesium II absorption lines in the QSOs spectra. This can be used to confirm that star formation is taking place. Part of one observation, and all of an other observation were lost because the Science Data Formatter was inadvertently disabled. 1.2 Completed Two Sets of WFC 4277 (Calibration fo Nearby Galaxy Type Ia Supernovae as Standard Candles) The Wide Field Camera was used to make a total of three observations of Cepheid variable stars located inside fo galaxy NGC-5253. This is to help determine the accurate distance scale to this galaxy. The distance will be used to find the true luminosity of the type - Ia supernova which has been observed in this galaxy. All three observations were reported to be satisfactory. 1.3 Completed FGS 4031 (High Speed Astrometry - A Search for Planetary Companions to Low Mass) The Fine Guidance Sensors were used to make astrometric observations of the star Proxima Centauri to look for the presence of a Jupiter sized planet in orbit around the target star based on the position perturbations caused by the planet. The OSS Monitored the observations as possible, and no problems were reported. 1.4 Completed PC 3603 (Snapshot of Protoplanetary Nebulae) The Planetary Camera was used to make snapshot images of the protoplanetary system M1-92 to study its current morphology. Two Sets of observations were received and displayed. One observation was monitored in the OSS, and was reported to be okay. Two other observations occurred during LOS, and are on the science tape recorders. They will be analyzed later. 1.5 Completed Two Sets of GHRS/2 4067 (GHRS 1ST Order Grating spectral Lamp Test) The High Resolution Spectrograph was used to make lamp mini-function wavelenght calibration observations for the first order gratings. This was strictly an internal test. No external sources were used. The observations were executed as scheduled, and no problems were reported. 1.6 Completed Two Set of FOC/96 3801 (Search for QSOs Suitable for Subsequent Observation of He II) The Faint Object Camera was used to attempt to make observations of quasar QSO 0043-259 to measure the levels of its ultraviolet flux, to see if the flux levels are high enough for subsequent observations using the GHRS and FOS. These observations are currently on the science tape recorder and will be played back later. No problems have been reported. 1.7 Completed PC 4712 (WF/PC Cycle 2 Calibration: Measles Monitor) The Planetary Camera was used to make Earth Calibration Flat images to monitor fo a form of contamination known as measles. The observations used filter F517N, both with and without filter F122M. The OSS monitored the observation and assessed the data quality. No problems were reported. 1.8 Completed WFC/PC 4109 (WF/PC Cycle 2 Calibration: Rapid Internal Monitor) Rapid Internal Monitoring observations were made with both the Wide Field Camera and the Planetary Camera using filter F555W to look for new forms of contamination (i.e. daises and measles), and the possible return of quantum efficiency hysteresis. The observations were monitored in the OSS, and the data quality was assessed. No problems were reported. 1.9 Completed Three Sets of FOC/96 4804 (UV Imaging of Nearby Galaxies With FOC/96) The Faint Object Camera was used to make six ultraviolet snapshot images of the nearby galaxies ESO-0220-2, UGC-1256, ESO-0044-2, ESO-0255-5, ESO-0306-2, and ESO-0011-2. Each exposure lasted ten minutes, and used filter F220W. All six images were received and displayed in the OSS. They all were reported to look okay. 1.10 Completed GHRS/2 4926 (GHRS SSA Pickup Test - Part 2) The High Resolution Spectrograph was used to make a calibration of the Small Science Aperture after a SAA peakup had been used to center the target. The peakup in the SAA failed, because the proper commanding has not been installed. The next three observations were received and displayed. These observations were made with the LSA and were reported to look okay. 1.11 Uplinked Updated Ephemeris Tables, Per MOSES 2871 1.12 Completed FOS/RD 4213 (Spectrophotometry of the Lensed "Clover Leaf" Broad Absorption line) The Faint Object Camera was used to make four observations of the Lensed quasi-stellar object QSO 1413+117 (aka, the Clover Leaf). A corrective maneuver was successfully calculated and uplinked to the spacecraft. The observations were displayed in the OSS, and they were reported to look okay. 1.13 Completed HSP/UV2 4209 (HSP Detector Sensitivity Test) The High Speed Photometer was used to make four calibration observations of standard star VID998 to determine the instruments long term sensitivity of the IDT detectors. All four observations were received and displayed. The first two observations showed the image of the target slightly off center. The next set of observations left it well centered. The following calibration images were reported to be good. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 20 Successful : 20 Scheduled Reacquisitions : 22 Successful : 22 Losses of Lock : 0 2.2 FHST Updates: Scheduled : 35 Successful : 35 2.3 Significant HSTARs: None 2.4 Operations Note: Ground Systems -------------- The transition from the PORTS system to the PRS system is under way in the STOCC. A successful Flight Readiness Review was conducted for this major installation last week. The PRS system will be run in "shadow" mode for several days as a final check-out. Benign ("no-op") commanding to the HST will be done using PRS after a successful shadow period. The PRS system will be brought on-line as the primary operational system in the MOR following a successful transition period. Fine Guidance Sensors --------------------- FGS 2 was used for guidance on June 10 for the first time since it was shut down. The acquisition was successful. There are five FGS 2 acquisitions scheduled during the next two weeks. Recommissioning of FGS 2 will continue over the next five weeks. At the end of this period the Project will assess the feasibility of returning FGS 2 to normal operations. 3.0 ACTIVITIES SCHEDULED FOR NEXT REPORTING PERIOD: 3.1 Complete WFC 4750 (WFC Observations of NGC-4319 - Markarian 205: High Resolution Morphology) 3.2 Complete FOC/96 3792 (UV Morphology of Starbursts in Interacting Galaxies) 3.3 Complete WFC 4381 (HST Medium - Deep Survey, Cycle 3) 3.4 Complete Two Sets of WFC 4713 (WF/PC Cycle 2 Calibration: WFC Internal Delta Flats) 3.5 Complete PC 4140 (WF/PC Cycle 2 Calibration: Rapid Internal Monitor - Part 2) 3.6 Complete FOC/96 4714 (WF/PC Cycle 2 Calibration: PC Internal Delta Flats) 3.7 Complete FOC/96 4846 (High Resolution Observations of Cataclysmic Variables: Cycle 3 High) 3.8 Complete WFC 4848 (GTO Low Latitude Parallel Program Cycle 3 High) 3.9 Uplink Updated Ephemeris Tables, Per MOSES 2871 3.10 Complete Two Sets of PC 3603 (Snapshot of Protoplanetary Nebulae) 3.11 Complete WFC/PC 4109 (WF/PC Cycle 2 Calibration: Rapid Internal Monitor) 3.12 Complete GHRS/2 4580 (High Resolution Observations of the Jovian UV Aurora _ Cycle 3 Med) 4.0 ANOMALY STATUS: 4.1 Significant New Anomalies: None 4.2 Significant Unresolved Anomalies: None 5.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. /DMH -END-